# A unique $Q$-point and infinitely many near-coherence classes of ultrafilters

### v2

@@ 1 Introduction
Throughout this paper, read ultrafilter as non-principal ultrafilter on $\omega$. For $x\subseteq\omega$, we denote by $[x]^{\omega}$ the set of infinite subsets of $x$ and by $\operatorname{fin}(x)$ the set of finite subsets of $x$.
@@
While the consistency of the non-existence of $Q$-points is a well-established fact with a variety of witnesses apart from the Mathias model333such as the Laver and Miller models (see [7] and [11], respectively)., the construction of models containing only ‘few’ $Q$-points seems to have received less attention. However, such models do arise naturally as models containing only few near-coherence classes of ultrafilters [^4] Two ultrafilters $\mathcal{U}_{1}$ and $\mathcal{U}_{2}$ are nearly-coherent iff there is some finite-to-one $f\in{{}^{\omega}}{\omega}$ such that $f(\mathcal{U}_{1})=f(\mathcal{U}_{2})$, where $f(\mathcal{U}_{i}):=\{X\subseteq\omega:f^{-1}[X]\in\mathcal{U}_{i}\}$. Note that two $Q$-points are nearly-coherent iff they are isomorphic.: Indeed, Mildenberger [10] has constructed models with exactly two and exactly three near-coherence classes, and it is easy to see that these contain exactly one and exactly two $Q$-points, respectively: In her model with exactly two near-coherence classes, one class contains a Ramsey ultrafilter, while the other class contains an ultrafilter that is $\omega_{1}$-generated. Hence, this latter class cannot contain a $Q$-point, since her models satisfy $\mathfrak{d}=\mathfrak{c}=\omega_{2}$ and such a $Q$-point would thus have to be ${<}\mathfrak{d}$-generated, which is impossible. Analogously, in Mildenberger’s model with exactly three near-coherence classes, two classes are represented by Ramsey ultrafilters, while the third contains an $\omega_{1}$-generated ultrafilter – giving exactly two $Q$-points in total.
@@
The construction of models with exactly $n$ near-coherence classes of ultrafilters for various finite $n\geq 4$ would similarly yield the consistency of exactly $m$ $Q$-points for some $m<n$ [^5] The inequality is strict since such a model must satisfy $\mathfrak{u}<\mathfrak{d}$, a result due to Banakh and Blass [9]. Hence, one of the $n$ near-coherence classes contains a ${<}\mathfrak{d}$-generated ultrafilter and thus no $Q$-point.
@@
The model considered in this paper is of a different nature, however: It contains only one $Q$-point while its number of near-coherence classes is $2^{\mathfrak{c}}$, i.e., the model’s lack of $Q$-points is not the consequence of a lack of near-coherence classes. This follows from the fact that dominating reals are added at each of the $\omega_{2}$-many stages of the iteration, which gives $\mathfrak{b}=\mathfrak{d}=\mathfrak{c}=\omega_{2}$ in the final extension. Since $\mathfrak{b}\leq\mathfrak{u}$ (see Solomon [8]), we have $\mathfrak{u}=\mathfrak{d}=\omega_{2}$ in our model, and hence there are $2^{\mathfrak{c}}$-many near-coherence classes of ultrafilters by Banakh and Blass [9].
@@
Let $\mathcal{U}$ be a Ramsey ultrafilter. Mathias forcing restricted to $\mathcal{U}$, written $\mathbb{M}_{\mathcal{U}}$, consists of conditions $\langle s,x\rangle\in\text{fin}(\omega)\times\mathcal{U}$ with $\max s<\min x$, ordered by
$\langle s,x\rangle\leq_{\mathbb{M}_{\mathcal{U}}}\langle t,y\rangle:\iff s\subseteq t\land x\supseteq y\land t\setminus s\subseteq x.$
Note that we use the convention that stronger forcing conditions are larger. The forcing notion $\mathbb{M}_{\mathcal{U}}$ clearly satisfies the c.c.c. and is therefore proper. We will need the following additional facts.
Fact 1.3 (e.g., see [1, Theorem 26.3]).
@@
Recall that a forcing notion $\mathbb{P}$ has the Laver property iff for every $\mathbb{P}$-name $\undertilde{g}$ for an element of ${{}^{\omega}}\omega$ such that there exists $f\in{{}^{\omega}}\omega\cap\mathbf{V}$ with
@@
we have that $\mathbb{P}$ forces that there exists $c:\omega\to\text{fin}(\omega)$ in $\mathbf{V}$ with
@@ 2 Result
Proof.
Assume that the ground model $\mathbf{V}$ satisfies CH. By induction, we define:
@@
    A countable support iteration $\mathbb{P}_{\omega_{2}}:=\langle\mathbb{P}_{\xi},,\undertilde{Q}_{\xi}:\xi\in\omega_{2}\rangle$ of c.c.c. forcing notions,
@@
Assume that we are in step $\xi\in\omega_{2}$. Let $G_{\xi}$ be $\mathbb{P}_{\xi}$-generic over $\mathbf{V}$ and work in $\mathbf{V}[G_{\xi}]$. Note that since $\mathbb{P}_{\xi}$ is a countable support iteration of proper forcing notions that are forced to be of size $\leq\omega_{1}$, we have $\mathbf{V}[G_{\xi}]\models\textsf{CH}$ (e.g., see [3, Theorem 2.12]). For each $\iota\in\xi$, let $\eta_{\iota}$ be the Mathias real added at stage $\iota$.
@@
Fact 2.1 (e.g., see  [3, Theorem 2.10]).
$\mathbb{P}_{\omega_{2}}$ is proper and satisfies the $\omega_{2}$-c.c..
We need to show that $\mathcal{U}_{\omega_{2}}$ is the only $Q$-point in $\mathbf{V}[G]$. To see this, assume by contradiction that $\mathbf{V}[G]\models``E\text{ is a }Q\text{-point and not isomorphic to }\mathcal{U}_{\omega_{2}}"$.
@@
There exists $\delta\in\omega_{2}$ such that $E\cap\mathbf{V}[G_{\delta}]\in\mathbf{V}[G_{\delta}]$ and $\mathbf{V}[G_{\delta}]\models``E\cap\mathbf{V}[G_{\delta}]\text{ is a }Q\text{-point and not isomorphic to }\mathcal{U}_{\delta}"$.
@@
    ``$\undertilde{e}_{\xi}$ is an enumeration (in $\omega_{1}$) of $\undertilde{E}\cap\mathbf{V}[\underaccent{\dot}{G}_{\xi}]"$. For each $\alpha\in\omega_{1}$ and $n\in\omega$ let $\mathcal{E}_{\xi,\alpha,n}\subseteq\mathbb{P}_{\omega_{2}}$ be a maximal antichain deciding ``$n\in\undertilde{e}_{\xi}(\alpha)"$.
@@
    ``$\undertilde{i}_{\xi}$ is an enumeration (in $\omega_{1}$) of the set of interval partitions of $\omega$ in $\mathbf{V}[\underaccent{\dot}{G}_{\xi}]"$. Note that we may assume that $\undertilde{i}_{\xi}$ is a $\mathbb{P}_{\xi}$-name.
@@
    “For all $\alpha\in\omega_{1}$, $\undertilde{s}_{\xi}(\alpha)$ is an element of $\undertilde{E}$ that intersects each interval in the interval partition $\undertilde{i}_{\xi}(\alpha)$ in at most one point”. Let $\mathcal{S}_{\xi,\alpha,n}\subseteq\mathbb{P}_{\omega_{2}}$ be a maximal antichain deciding ``$n\in\undertilde{s}_{\xi}(\alpha)"$.
@@
    ``$\undertilde{b}_{\xi}$ is an enumeration (in $\omega_{1}$) of all permutations of $\omega$ in $\mathbf{V}[\underaccent{\dot}{G}_{\xi}]"$. We may again assume that $\undertilde{b}_{\xi}$ is a $\mathbb{P}_{\xi}$-name.
@@
    “For all $\alpha\in\omega_{1}$, $\undertilde{f}_{\xi}(\alpha)$ is a pair $\text{op}(\undertilde{x}_{\alpha},\undertilde{y}_{\alpha})$ such that $\undertilde{x}_{\alpha}$ is in $\undertilde{E}$, $\undertilde{y}_{\alpha}$ is in $\undertilde{\mathcal{U}}_{\omega_{2}}$ and $\undertilde{b}_{\xi}(\alpha)[\undertilde{x}_{\alpha}]$ is disjoint from $\undertilde{y}_{\alpha}"$. Let $\mathcal{X}_{\xi,\alpha,n}\subseteq\mathbb{P}_{\omega_{2}}$ be a maximal antichain deciding ``$n\in\undertilde{x}_{\alpha}"$, and define $\mathcal{Y}_{\xi,\alpha,n}$ analogously.
By the $\omega_{2}$-c.c. of $\mathbb{P}_{\omega_{2}}$, there exists for each $\xi\in\omega_{2}$ some $\gamma_{\xi}\in\omega_{2}$ greater than $\xi$ such that all the above antichains consist of $\mathbb{P}_{\gamma_{\xi}}$-conditions. Recursively define $\lambda(0)=0$, $\lambda(\xi+1)=\gamma_{\lambda(\xi)}$ and for limit ordinals $\xi:\lambda(\xi)=\bigcup_{\iota\in\xi}\lambda(\iota)$, for $\xi\leq\omega_{1}$. Set $\delta:=\lambda(\omega_{1})$ and consider the extension $\mathbf{V}[G_{\delta}]$. Since $\text{cf}(\delta)=\omega_{1}$, we have that $E\cap\mathbf{V}[G_{\delta}]=\bigcup_{\iota\in\omega_{1}}E\cap\mathbf{V}[G_{\lambda(\iota)}]$, and since each $E\cap\mathbf{V}[G_{\lambda(\iota)}]$ is an element of $\mathbf{V}[G_{\delta}]$ by (i), $E\cap\mathbf{V}[G_{\delta}]$ is an element of $\mathbf{V}[G_{\delta}]$ (and an ultrafilter). Furthermore, any interval partition of $\omega$ in $\mathbf{V}[G_{\delta}]$ already appears in some $\mathbf{V}[G_{\lambda(\iota)}]$, $\iota\in\omega_{1}$, where it equals $\undertilde{i}_{\lambda(\iota)}[G_{\lambda(\iota)}](\alpha)$ for some $\alpha\in\omega_{1}$. Since $\undertilde{s}_{\lambda(\iota)}[G_{\delta}](\alpha)\in E\cap\mathbf{V}[G_{\delta}]$, we obtain that $E\cap\mathbf{V}[G_{\delta}]$ is a $Q$-point. Finally and analogously, any permutation of $\omega$ in $\mathbf{V}[G_{\delta}]$ already appears in $\mathbf{V}[G_{\lambda(\iota)}]$ for some $\iota\in\omega_{1}$ and hence there are witnesses $\undertilde{x}_{\alpha}[G_{\delta}]\in E\cap\mathbf{V}[G_{\delta}]$ and $\undertilde{y}_{\alpha}[G_{\delta}]\in\mathcal{U}_{\omega_{2}}\cap\mathbf{V}[G_{\delta}]=\mathcal{U}_{\delta}$ witnessing that $E\cap\mathbf{V}[G_{\delta}]$ and $\mathcal{U}_{\delta}$ are not isomorphic.   $\dashv$
We now designate $\mathbf{V}[G_{\delta}]$ as the new ground model and rename the $Q$-point $E\cap\mathbf{V}[G_{\delta}]$ to $E$ and the Ramsey ultrafilter $\mathcal{U}_{\delta}$ to $\mathcal{U}$. Note that by the Factor-Lemma (e.g., see [4, Theorem 4.6]), the quotient $\mathbb{P}_{\omega_{2}}/G_{\delta}$ is again isomorphic to a countable support iteration of restricted Mathias forcings. In particular, by Facts 1.5 and 1.6, $\mathbb{P}_{\omega_{2}}/G_{\delta}$ is isomorphic to the two-step iteration $\mathbb{M}_{\mathcal{U}}\ast\undertilde{R}$, where $\mathbb{M}_{\mathcal{U}}\Vdash``\undertilde{R}\text{ has the Laver property}"$.
It remains to show the following.
@@
then there exists some $v\in E$ and some $\langle\bar{p},\undertilde{\bar{q}}\rangle$ greater than $\langle p,\undertilde{q}\rangle$ such that
@@
Note that $\undertilde{a}$ is forced by $\mathbb{M}_{\mathcal{U}}$ to be dominated by $\underaccent{\dot}{\eta}$. Hence, by the Laver property of $\undertilde{R}$, there exists a $\mathbb{M}_{\mathcal{U}}$-name $\undertilde{c}$ for a function from $\omega$ to $\text{fin}(\omega)$ and some $\langle p^{\prime},\undertilde{q}^{\prime}\rangle\geq_{\mathbb{M}_{\mathcal{U}}\ast\undertilde{R}}\langle p,\undertilde{q}\rangle$ such that
@@
We may assume without loss of generality that $p^{\prime}\Vdash_{\mathbb{M}_{\mathcal{U}}}\forall n\in\omega:\undertilde{c}(n)\subseteq(\underaccent{\dot}{\eta}(n-1),\underaccent{\dot}{\eta}(n)]$. Let $\undertilde{C}$ be a $\mathbb{M}_{\mathcal{U}}$-name for an element of $[\omega]^{\omega}$ such that $p^{\prime}\Vdash_{\mathbb{M}_{\mathcal{U}}}\undertilde{C}=\bigcup\text{range}(\undertilde{c})$. Hence, we have
@@
Write $p^{\prime}=\langle s,x_{0}\rangle$. There exists $x_{1}\in[x_{0}]^{\omega}\cap\mathcal{U}$ such that the $\mathbb{M}_{\mathcal{U}}$-condition $\langle s,x_{1}\rangle\geq_{\mathbb{M}_{\mathcal{U}}}\langle s,x_{0}\rangle$ has the following property:
For every $t\in\text{fin}(x_{1})$, there exists $C_{t}\in\text{fin}(\omega)$ such that
@@
For every $t\in\text{fin}(x_{2})$, every $m\in x_{2}\setminus\max t$ and all $n,n^{\prime}\in x_{2}\setminus m^{+}$, it holds that $C_{t\cup\{n\}}\cap m^{+}=C_{t\cup\{n^{\prime}\}}\cap m^{+}$.
@@
Note that the conclusion of Lemmas 2.4 and 2.5 also holds for each $\langle s^{\prime},x^{\prime}\rangle\geq_{\mathbb{M}_{\mathcal{U}}}\langle s,x_{2}\rangle$. Hence, we simply trim $x_{2}$ such that the enumeration of $s\cup x_{2}$ dominates $2^{j+1}$ above $|s|$ and replace $s$ with $s\cup\{\min x_{2}\}$ and $x_{2}$ with $x_{2}\setminus\{\min x_{2}\}$.   $\dashv$
@@
    $\forall i\in\omega:\forall t\in\text{fin}(\omega):t\in N_{i}\iff t\subseteq N_{i}$,
@@
    If $\langle m,l,D\rangle\in(\omega\times\omega\times\text{fin}(\omega))\cap N_{i}$, then $m,l,D\in N_{i}$ (and hence $D\subseteq N_{i}$ by the previous condition).
@@
Assume $\langle s\cup t,x^{\prime}\rangle\geq_{\mathbb{M}_{\mathcal{U}}}\langle s,x_{2}\rangle$, $a\in\omega\setminus(\max s)^{+}$ and $i\in\omega\setminus\{0,1\}$ are such that
@@
If $l^{*}\in[k_{i-1},k_{i})$, we are done, since this means that $\langle s\cup t,x^{\prime}\rangle\Vdash_{\mathbb{M}_{\mathcal{U}}}l^{*}\in\text{range}(\underaccent{\dot}{\eta})\cap[k_{i-1},k_{i})$. Hence, assume $l^{*}\notin[k_{i-1},k_{i})$, which means that $l^{*}\notin N_{i}$, since $l^{*}$ is certainly not in $N_{i-1}$ (if it were, $a$ would be as well by (iii)). Note that $l^{*}$ witnesses that
@@
Hence, by (v), we have that $l^{*}\in N_{i+1}$ and thus $l^{*}\in[k_{i},k_{i+1})$.
@@
Let $j\in\omega$ be such that $l^{*}$ is the $j$’th elements of $s\cup t$ in increasing order. By Lemmas 2.4 and 2.5, there is $C_{t^{\prime}\cup\{l^{*}\}}\subseteq(l^{*})^{+}$ such that
@@
$\langle s\cup t^{\prime}\cup\{l^{*}\},x_{2}\setminus(l^{*})^{+}\rangle\leq_{\mathbb{M}_{\mathcal{U}}}\langle s\cup t,x^{\prime}\rangle,$
@@
Since $l_{0}$ is the $(j-1)$’th element of $s\cup t^{\prime}$, we have $l_{0}>2^{j}$ by Fact 2.6.666Note that the additional requirement in Fact 2.6 that $\max s$ is already larger than $2^{|s|}$ is needed here, since $l_{0}$ could be $\max s$. Hence, since $l_{0}\in N_{i^{*}}$, it follows that $\gamma\in N_{i^{*}}$. Thus, all the parameters in the above formula lie in $N_{i^{*}}$, which implies that there exists $\langle m^{\dagger},l^{\dagger},D^{\dagger}\rangle\in N_{i^{*}+1}$ satisfying the formula.
@@
Proof. Assume by contradiction that $l^{\dagger}<a$, i.e.,

### v3

# arXiv:2505.17960v3[math.LO] 01 Mar 2026

@@ 0 Introduction
Throughout this paper, read ultrafilter as non-principal ultrafilter on $\omega$. For $x\subseteq\omega$, we denote by $x]^{\omega}$ the set of infinite subsets of $x$ and by $x]^{<\omega}$ the set of finite subsets of $x$.
@@
It is well-known that every Ramsey ultrafilter is a $Q$-point. Canjar [4] showed that the existence of $2^{\mathfrak{c}}$-many Ramsey ultrafilters follows from the assumption $\mathit{cov}(\mathcal{M})=\mathfrak{c}$. The weaker assumption $\mathit{cov}(\mathcal{M})=\mathfrak{d}$ implies the existence of $2^{\mathfrak{c}}$ $Q$-points, as was shown by Millán [12]. It is well-known that in the Mathias model – the model obtained by a length-$\omega_{2}$ countable support iteration of unrestricted Mathias forcing over a ground model satisfying CH – there are no $Q$-points (see [9, Proposition 26.23]). In fact, the Mathias model contains no rapid ultrafilters, where an ultrafilter $E$ is rapid if for every $f\in{}^{\omega}\omega$ there exists some $x\in E$ such that $\forall n\in\omega:|x\cap f(n)|\leq n$ (note that every $Q$-point is rapid). It follows that both the Mathias model and the model considered in this paper satisfy $\mathit{cov}(\mathcal{M})=\omega_{1}<\mathfrak{d}=\mathfrak{c}=\omega_{2}$.
@@
While the consistency of the non-existence of $Q$-points is a well-established fact with a variety of witnesses apart from the Mathias model333such as the Laver and Miller models (see [13] and [3], respectively)., the construction of models containing only ‘few’ $Q$-points seems to have received less attention. However, such models do arise naturally as models containing only few near-coherence classes of ultrafilters [^4] Two ultrafilters $\mathcal{U}_{1}$ and $\mathcal{U}_{2}$ are nearly-coherent if there is some finite-to-one $f\in{{}^{\omega}}{\omega}$ such that $f(\mathcal{U}_{1})=f(\mathcal{U}_{2})$, where $f(\mathcal{U}_{i}):=\{X\subseteq\omega:f^{-1}[X]\in\mathcal{U}_{i}\}$. Note that two $Q$-points are nearly-coherent if and only if they are isomorphic.: Indeed, Mildenberger [11] has constructed models with exactly two and exactly three near-coherence classes, and it is easy to see that these contain exactly one and exactly two $Q$-points, respectively: In her model with exactly two near-coherence classes, one class contains a Ramsey ultrafilter, while the other class contains an ultrafilter that is $\omega_{1}$-generated. Hence, this latter class cannot contain a $Q$-point, since her models satisfy $\mathfrak{d}=\mathfrak{c}=\omega_{2}$ and such a $Q$-point would thus have to be $\mathfrak{d}-$generated, which is impossible. Analogously, in Mildenberger’s model with exactly three near-coherence classes, two classes are represented by Ramsey ultrafilters, while the third contains an $\omega_{1}$-generated ultrafilter – giving exactly two $Q$-points in total.
@@
The construction of models with exactly $n$ near-coherence classes of ultrafilters for various finite $n\geq 4$ would similarly yield the consistency of exactly $m$ $Q$-points for some m<nm<n. [^5] The inequality is strict since such a model must satisfy $\mathfrak{u}<\mathfrak{d}$, a result due to Banakh and Blass [2]. Hence, one of the $n$ near-coherence classes contains a $\mathfrak{d}-$generated ultrafilter and thus no $Q$-point. [^6] See the note on the next page.
@@
The model considered in this paper is of a different nature, however: It contains only one $Q$-point while its number of near-coherence classes is $2^{\mathfrak{c}}$, i.e., the model’s lack of $Q$-points is not the consequence of a lack of near-coherence classes. This follows from the fact that dominating reals are added at each of the $\omega_{2}$ stages of the iteration, which gives $\mathfrak{b}=\mathfrak{d}=\mathfrak{c}=\omega_{2}$ in the final extension. Since $\mathfrak{b}\leq\mathfrak{u}$ (see Solomon [15]), we have $\mathfrak{u}=\mathfrak{d}=\omega_{2}$ in our model, and hence there are $2^{\mathfrak{c}}$-many near-coherence classes of ultrafilters by Banakh and Blass [2].
Incidentally, both our model as well as Mildenberger’s models answer the two questions posed in a recent paper of Raghavan [5]. He asked whether it is consistent that there are $Q$-points, while (1) there is no Tukey maximal $Q$-point, and (2) the $Q$-points are not cofinal in the RK-ordering. [^7] See, for example, [6] and [7] on the Tukey- and RK-orderings of ultrafilters. Note that the unique $Q$-point in our model is simultaneously a $P$-point, and these are never Tukey maximal (see [7, Corollary 19]).
Note added in proof:
Based on the technique used below, in the forthcoming paper [10], Halbeisen, Horvath, and Özalp have established the consistency of the statement “There are exactly $n$ $Q$-points up to isomorphism and $\mathfrak{u}=\mathfrak{d}=\mathfrak{c}$, for any $n\in\omega$.
1 Definitions and Preliminaries
Before we state and prove our main result in the next section, let us give some definitions, relations, and related results.
@@ 1 Definitions and Preliminaries
Let $\mathcal{U}$ be a Ramsey ultrafilter. Mathias forcing restricted to $\mathcal{U}$, written $\mathbb{M}_{\mathcal{U}}$, consists of conditions $\langle s,x\rangle\in[\omega]^{<\omega}\times\mathcal{U}$ with $\max s<\min x$, ordered by
$\langle s,x\rangle\leq_{\mathbb{M}_{\mathcal{U}}}\langle t,y\rangle:\iff s\supseteq t\land x\subseteq y\land s\setminus t\subseteq y.$
The forcing notion $\mathbb{M}_{\mathcal{U}}$ clearly satisfies the c.c.c. and is therefore proper. We will need the following additional facts.
Fact 1.2 (e.g., see [9, Theorem 26.3]).
@@
Recall that a forcing notion $\mathbb{P}$ has the Laver property if for every $\mathbb{P}$-name $\undertilde{g}$ for an element of ${{}^{\omega}}\omega$ such that there exists $f\in{{}^{\omega}}\omega\cap\mathbf{V}$ with
@@
we have that $\mathbb{P}$ forces that there exists $c:\omega\to[\omega]^{<\omega}$ in $\mathbf{V}$ with
@@ 2 Result
The proof is given by the following construction and the subsequent results: Assume that the ground model $\mathbf{V}$ satisfies CH. By induction, we define:
@@
    A countable support iteration $\mathbb{P}_{\omega_{2}}:=\langle\mathbb{P}_{\xi},\undertilde{Q}_{\xi}:\xi\in\omega_{2}\rangle$ of c.c.c. forcing notions,
@@
Assume that we are in step $\xi\in\omega_{2}$. Let $G_{\xi}$ be $\mathbb{P}_{\xi}$-generic over $\mathbf{V}$ and work in $\mathbf{V}[G_{\xi}]$. Note that since $\mathbb{P}_{\xi}$ is a countable support iteration of proper forcing notions that are forced to be of size $\leq\omega_{1}$, we have $\mathbf{V}[G_{\xi}]\models\textsf{CH}$ (e.g., see [1, Theorem 2.12]). For each $\iota\in\xi$, let $\eta_{\iota}$ be the Mathias real added at stage $\iota$.
@@
Fact 2.1 (see, e.g., [1, Theorem 2.10]).
$\mathbb{P}_{\omega_{2}}$ is proper and satisfies $\omega_{2}$-$c.$.
[…]
@@
[…]
@@
    $\undertilde{e}_{\xi}$ is an enumeration (in $\omega_{1}$) of $\undertilde{E}\cap\mathbf{V}[\underaccent{\dot}{G}_{\xi}]$". For each $\alpha\in\omega_{1}$ and $n\in\omega$ let $\mathcal{E}_{\xi,\alpha,n}\subseteq\mathbb{P}_{\omega_{2}}$ be a maximal antichain deciding $n\in\undertilde{e}_{\xi}(\alpha)$".
@@
    $\undertilde{i}_{\xi}$ is an enumeration (in $\omega_{1}$) of the set of interval partitions of $\omega$ in $\mathbf{V}[\underaccent{\dot}{G}_{\xi}]$". Note that we may assume that $\undertilde{i}_{\xi}$ is a $\mathbb{P}_{\xi}$-name.
@@
    “For all $\alpha\in\omega_{1}$, $\undertilde{s}_{\xi}(\alpha)$ is an element of $\undertilde{E}$ that intersects each interval in the interval partition $\undertilde{i}_{\xi}(\alpha)$ in at most one point”. Let $\mathcal{S}_{\xi,\alpha,n}\subseteq\mathbb{P}_{\omega_{2}}$ be a maximal antichain deciding $n\in\undertilde{s}_{\xi}(\alpha)$".
@@
    $\undertilde{b}_{\xi}$ is an enumeration (in $\omega_{1}$) of all permutations of $\omega$ in $\mathbf{V}[\underaccent{\dot}{G}_{\xi}]$". We may again assume that $\undertilde{b}_{\xi}$ is a $\mathbb{P}_{\xi}$-name.
@@
    “For all $\alpha\in\omega_{1}$, $\undertilde{f}_{\xi}(\alpha)$ is a pair $\langle\undertilde{x}_{\alpha},\undertilde{y}_{\alpha}\rangle$ such that $\undertilde{x}_{\alpha}$ is in $\undertilde{E}$, $\undertilde{y}_{\alpha}$ is in $\undertilde{\mathcal{U}}_{\omega_{2}}$ and $\undertilde{b}_{\xi}(\alpha)[\undertilde{x}_{\alpha}]$ is disjoint from $\undertilde{y}_{\alpha}$". Let $\mathcal{X}_{\xi,\alpha,n}\subseteq\mathbb{P}_{\omega_{2}}$ be a maximal antichain deciding $n\in\undertilde{x}_{\alpha}$", and define $\mathcal{Y}_{\xi,\alpha,n}$ analogously.
Since $\mathbb{P}_{\omega_{2}}$ satisfies $\omega_{2}$-c.c., there exists for each $\xi\in\omega_{2}$ some $\gamma_{\xi}\in\omega_{2}$ greater than $\xi$ such that all the above antichains consist of $\mathbb{P}_{\gamma_{\xi}}$-conditions. Recursively define $\lambda(0)=0$, $\lambda(\xi+1)=\gamma_{\lambda(\xi)}$ and for limit ordinals $\xi:\lambda(\xi)=\bigcup_{\iota\in\xi}\lambda(\iota)$, for $\xi\leq\omega_{1}$. Set $\delta:=\lambda(\omega_{1})$ and consider the extension $\mathbf{V}[G_{\delta}]$. Since $\text{cf}(\delta)=\omega_{1}$, we have that $E\cap\mathbf{V}[G_{\delta}]=\bigcup_{\iota\in\omega_{1}}E\cap\mathbf{V}[G_{\lambda(\iota)}]$, and since each $E\cap\mathbf{V}[G_{\lambda(\iota)}]$ is an element of $\mathbf{V}[G_{\delta}]$ by (i), $E\cap\mathbf{V}[G_{\delta}]$ is an element of $\mathbf{V}[G_{\delta}]$ (and an ultrafilter). Furthermore, any interval partition of $\omega$ in $\mathbf{V}[G_{\delta}]$ already appears in some $\mathbf{V}[G_{\lambda(\iota)}]$, $\iota\in\omega_{1}$, where it equals $\undertilde{i}_{\lambda(\iota)}[G_{\lambda(\iota)}](\alpha)$ for some $\alpha\in\omega_{1}$. Since $\undertilde{s}_{\lambda(\iota)}[G_{\delta}](\alpha)\in E\cap\mathbf{V}[G_{\delta}]$, we obtain that $E\cap\mathbf{V}[G_{\delta}]$ is a $Q$-point. Finally and analogously, any permutation of $\omega$ in $\mathbf{V}[G_{\delta}]$ already appears in $\mathbf{V}[G_{\lambda(\iota)}]$ for some $\iota\in\omega_{1}$ and hence there are witnesses $\undertilde{x}_{\alpha}[G_{\delta}]\in E\cap\mathbf{V}[G_{\delta}]$ and $\undertilde{y}_{\alpha}[G_{\delta}]\in\mathcal{U}_{\omega_{2}}\cap\mathbf{V}[G_{\delta}]=\mathcal{U}_{\delta}$ witnessing that $E\cap\mathbf{V}[G_{\delta}]$ and $\mathcal{U}_{\delta}$ are not isomorphic. $\dashv$
We now designate $\mathbf{V}[G_{\delta}]$ as the new ground model and rename the $Q$-point $E\cap\mathbf{V}[G_{\delta}]$ to $E$ and the Ramsey ultrafilter $\mathcal{U}_{\delta}$ to $\mathcal{U}$. Note that by the Factor-Lemma (e.g., see [8, Theorem 4.6]), the quotient $\mathbb{P}_{\omega_{2}}/G_{\delta}$ is again isomorphic to a countable support iteration of restricted Mathias forcings. In particular, by Facts 1.4 and 1.5, $\mathbb{P}_{\omega_{2}}/G_{\delta}$ is isomorphic to the two-step iteration $\mathbb{M}_{\mathcal{U}}\ast\undertilde{R}$, where $\mathbb{M}_{\mathcal{U}}\Vdash``\undertilde{R}\text{ has the Laver property}$".
It remains to show the following
@@
then there exists some $v\in E$ and some $\langle\bar{p},\undertilde{\bar{q}}\rangle\leq_{\mathbb{M}_{\mathcal{U}}\ast\undertilde{R}}\langle p,\undertilde{q}\rangle$ such that
@@
Note that $\undertilde{a}$ is forced by $\mathbb{M}_{\mathcal{U}}$ to be dominated by $\underaccent{\dot}{\eta}$. Hence, by the Laver property of $\undertilde{R}$, there exists a $\mathbb{M}_{\mathcal{U}}$-name $\undertilde{c}$ for a function from $\omega$ to $\omega]^{<\omega}$ and some $\langle p^{\prime},\undertilde{q}^{\prime}\rangle\leq_{\mathbb{M}_{\mathcal{U}}\ast\undertilde{R}}\langle p,\undertilde{q}\rangle$ such that
@@
[…]
@@
Write $p^{\prime}=\langle s,x_{0}\rangle$. There exists $x_{1}\in[x_{0}]^{\omega}\cap\mathcal{U}$ such that the $\mathbb{M}_{\mathcal{U}}$-condition $\langle s,x_{1}\rangle\leq_{\mathbb{M}_{\mathcal{U}}}\langle s,x_{0}\rangle$ has the following property:
For every $t\in[x_{1}]^{<\omega}$, there exists $C_{t}\in[\omega]^{<\omega}$ such that
@@
For every $t\in[x_{2}]^{<\omega}$, every $m\in x_{2}\setminus\max t$ and all $n,n^{\prime}\in x_{2}\setminus m^{+}$, it holds that $C_{t\cup\{n\}}\cap m^{+}=C_{t\cup\{n^{\prime}\}}\cap m^{+}$.
@@
Note that the conclusion of Lemmata 2.4 and 2.5 also holds for each $\langle s^{\prime},x^{\prime}\rangle\leq_{\mathbb{M}_{\mathcal{U}}}\langle s,x_{2}\rangle$. Hence, we simply trim $x_{2}$ such that the enumeration of $s\cup x_{2}$ dominates $2^{j+1}$ above $|s|$ and replace $s$ with $s\cup\{\min x_{2}\}$ and $x_{2}$ with $x_{2}\setminus\{\min x_{2}\}$. $\dashv$
@@
    $\forall i\in\omega:\forall t\in[\omega]^{<\omega}:t\in N_{i}\iff t\subseteq N_{i}$,
@@
    If $\langle m,l,D\rangle\in(\omega\times\omega\times[\omega]^{<\omega})\cap N_{i}$, then $m,l,D\in N_{i}$ (and hence $D\subseteq N_{i}$ by the previous condition).
@@
Assume $\langle s\cup t,x^{\prime}\rangle\leq_{\mathbb{M}_{\mathcal{U}}}\langle s,x_{2}\rangle$, $a\in\omega\setminus(\max s)^{+}$ and $i\in\omega\setminus\{0,1\}$ are such that
@@
If $l^{*}\in[k_{i-1},k_{i})$, we are done, since this means that $\langle s\cup t,x^{\prime}\rangle\Vdash_{\mathbb{M}_{\mathcal{U}}}l^{*}\in\text{range}(\underaccent{\dot}{\eta})\cap[k_{i-1},k_{i})$. Hence, assume $l^{*}\notin[k_{i-1},k_{i})$, i.e., $l^{*}\notin N_{i}$. Note that $l^{*}$ witnesses that
@@
Hence, by (vi), we have that $l^{*}\in N_{i+1}$ and thus $l^{*}\in[k_{i},k_{i+1})$.
@@
Let $j\in\omega$ be such that $l^{*}$ is the $j$’th elements of $s\cup t$ in increasing order. By Lemmata 2.4 and 2.5, there is $C_{t^{\prime}\cup\{l^{*}\}}\subseteq(l^{*})^{+}$ such that
@@
$\langle s\cup t^{\prime}\cup\{l^{*}\},x_{2}\setminus(l^{*})^{+}\rangle\geq_{\mathbb{M}_{\mathcal{U}}}\langle s\cup t,x^{\prime}\rangle,$
@@
Since $l_{0}$ is the (j1)(j-1)’th element of $s\cup t^{\prime}$, we have $l_{0}>2^{j}$ by Fact 2.6.888Note that the additional requirement in Fact 2.6 that $\max s$ is already larger than $2^{|s|}$ is needed here, since $l_{0}$ could be $\max s$. Hence, since $l_{0}\in N_{i^{*}}$, it follows that $\gamma\in N_{i^{*}}$. Thus, all the parameters in the above formula lie in $N_{i^{*}}$, which implies that there exists $\langle m^{\dagger},l^{\dagger},D^{\dagger}\rangle\in N_{i^{*}+1}$ satisfying the formula.
@@
Proof of Claim. Assume by contradiction that $l^{\dagger}<a$, i.e.,